A particle moves along the parabola y2 = 4x with a constant horizontal component velocity of 2 m/s. Find the vertical component of the velocity at the point (1, 2).


1
Expert's answer
2022-06-09T13:02:41-0400


We know that:

"v_x=\frac{dx}{dt}," "v_y=\frac{dy}{dt}," and given "|\vec{v_x}|=2" m/s.


Let's differentiate the function "y^2=4x" :

"2y\frac{dy}{dt}=4\frac{dx}{dt},"


"2yv_y=4v_x,"


"v_y=\frac{4v_x}{2y}=\frac{2v_x}{y}=\frac{2\cdot2}{y}=\frac{4}{y}."


At the point (1, 2) we have y=2, so

"v_y=\frac{4}{2}=2" m/s.


Answer: 2 m/s.